# Configuring Agentic AI Coding Tools: An Exploratory Study

### v1

@@ 5.1. Characterization of Repositories in Dataset
Before analyzing adoption, we characterize the 2,926 repositories in our dataset to contextualize the configuration mechanisms of agentic AI coding tools.
Most repositories adopted a single tool. Among repositories with multiple tools, Claude appeared most frequently with others. The most common combination was Claude and Copilot ($n=128$). Overlaps of three or more tools were rare. In our initial sample, the top five main programming languages per repository ($n=36,194$) were Python (8,133; 22.5%); TypeScript (4,999; 13.8%); Java (3,813; 10.5%); Go (3,786; 10.5%); and JavaScript (3,709; 10.3%). Interestingly, this order is slightly different for repositories that use agentic coding tools ($n=2,926$, see Figure 3) where TypeScript and Go were more prominent and C# replaces JavaScript: TypeScript (922, 31.5%); Python (548, 18.7%); Go (533, 18.2%); C# (256, 8.8%); Java (246, 8.4%). TypeScript is the most common primary language for repositories using Claude, Codex, Copilot, and Cursor. Usage in Java and C# repositories is lower for all tools but Copilot. As shown in Table 3, repositories using Cursor are younger compared to those using other tools. Repositories associated with Gemini exhibit larger contributor counts and commit volumes, compared to those associated with other tools. Cursor repositories are the largest in terms of kilobytes of source code.
@@ 5.2. Distribution of Configuration Mechanisms
Figure 3 shows the distribution of configuration mechanisms. Across repositories, Context Files (with artifacts such as CLAUDE.md or AGENTS.md files) were the most frequently adopted configuration mechanism. In contrast, other configuration mechanisms, including Rules, Settings, Commands, Skills, and Subagents, were substantially less prevalent. Each of these mechanisms appeared in fewer than 20% of the repositories. Rules are primarily concentrated in repositories using Cursor, since Cursor was one of the first to introduce Rules (Jiang and Nam, 2026).
Most repositories adopt a limited subset of available configuration mechanisms. In many cases, repositories include only a single Context File artifact. The adoption of multiple non-context file configuration mechanisms within the same repository remains comparatively rare. These results suggest that configuration usage is currently characterized by one dominant baseline configuration (i.e., Context Files).
Figure 4. Cumulative adoption of selected configuration artifacts for agentic tools. Copilot Instructions include the two different artifact types for Context Files, see Table 1.
@@
Diagram showing the comulative adoption of configuraiton artifacts for agentic tools.
@@ 5.2.1. Co-occurrence of Configuration Mechanisms
Several configuration mechanisms are frequently adopted together. Some of these correlations can be explained by shared configuration artifacts. For example, Settings and Hooks have a strong positive correlation ($\rho=0.52$), because hooks are usually defined in Settings files. However, Settings and Hooks also both co-occur with Skills ($\rho=0.23$ and $\rho=0.24$, respectively), which are defined and configured differently. Subagents also show positive correlations with multiple mechanisms, including Settings, Skills, and MCP. Other configuration mechanisms appear less frequently in combination. Context Files are negatively correlated with Rules ($\rho=-0.36$) and Commands ($\rho=-0.14$), which can again attributed to how they are defined and which tools first introduced the mechanisms. Overall, configuration mechanism adoption across repositories is characterized by recurring combinations of mechanisms rather than uniform usage. Patters of co-occurrence might be impacted by the tools and configuration mechanisms they support (e.g., Claude, the dominating tool, does not support Rules).
@@ 5.2.2. Adoption of Configuration Mechanisms over Time
In Figure 5, we show the adoption of configuration mechanisms over time.
The figure shows individual artifacts that together form the mechanism Context Files, plus Skills and Subagents which have a uniform format.
Context Files (CLAUDE.md, AGENTS.md, copilot-instructions.md) clearly dominate and increase continuously, while Skills and Subagents experienced a comparatively slow growth. As the figure shows, .cursorrules and copilot-instructions.md started being introduced in 2024. Although Cursor and also Copilot were originally released in 2023 and 2021, respectively, agentic capabilities were only introduced in 2025 (see Table 2). The adoption of copilot-insturctions.md has increased since then.
@@
Key insights for RQ2:
  • Context Files are the dominant and often sole configuration mechanism. Advanced configuration mechanisms (e.g., Hooks, MCP) are adopted less frequently.
  • Adoption varies: Claude has a broad configuration footprint, while Cursor emphasizes rule-based mechanisms.
  • Although certain correlations between configuration mechanisms exist, the landscape still appears in flux.
@@ 6. Details of Configuration Mechanisms (RQ3)
In this section, we analyze the use of three configuration mechanisms in more detail. We selected Context Files and Skills because they are supported by all tools, as discussed in Section 5, and Subagents because they follow a similar format as Skills.
@@ 6.1. Configuration Mechanism: Context Files
Context Files are Markdown files that provide a central machine-readable source of contextual information about the repository in which they are used. AGENTS.md was introduced by OpenAI and now serves as an open tool-agnostic convention that is increasingly supported by various tools (Mohsenimofidi et al., 2026). We analyzed the presence and co-occurrence of Context Files across repositories. We identified 4,860 Context Files across 2,634 of the 2,926 repositories in our sample. The histogram in Figure 5 shows the distribution of Context Files. Most repositories that do use Context Files include one or two of such files.
CLAUDE.md emerges as the dominant file type with 1,661 (34.2%), followed closely by AGENTS.md and copilot-instructions.md with 1,572 (32.3%) and 1,344 (27.7%) files, respectively. GEMINI.md (159 files, 3.3%) and .cursorrules (73 files, 1.5%) are rare. Note that .cursorrules are now deprecated; Cursor suggests using AGENTS.md instead. Cursor rules are different and are stored in .cursor/rules/. We also investigated the repository-level adoption of the Context Files, where CLAUDE.md demonstrates the highest adoption rate with 45.40% of repositories (1,195 repos), AGENTS.md follows closely with 40.6% adoption (1,069 repos), and copilot-instructions.md appears in 35.1% of repositories (925 repos). The predominance of CLAUDE.md, AGENTS.md and copilot-instructions.md, both in terms of their frequency and the number of repositories adopting them, indicates a clear concentration of projects around these files.
We then investigated programming language-specific patterns in Context Files usage across the repositories. We found Context Files in 651 repositories with TypeScript as the main language, 435 with Python, 414 with Go, 210 with Java, 210 with C#, 201 with JavaScript, 134 with Ruby, 181 with Rust, 156 with C++, 123 with PHP, and 53 with C. CLAUDE.md is the most common Context File across languages, except for Java, C#, and C++, where copilot-instructions.md emerges as the dominant type.
@@
Diagram showing the creation order of Context Files per repository.
Figure 7 shows the creation order of the different Context Files. Most repositories only used one type of Context Files. These repositories are not captured in this figure. AGENTS.md and CLAUDE.md are typically created first, and if CLAUDE.md existed first, then AGENTS.md was commonly created afterward. A reason for this could be that Claude Code is the most popular agentic AI coding tool, but does not yet support the emerging standard AGENTS.md (anthropics/claude-code on GitHub, 2026). Interestingly, repositories that started with copilot-instructions.md then often created CLAUDE.md or AGENTS.md files, even though GitHub Copilot supports all major Context Files types. This, in addition to the overall dominance of CLAUDE.md and AGENTS.md, illustrates that these two files currently serve as the de facto standard.
Finally, we observed that Context Files commonly refer to each other. This is visualized in Figure 7. “Referring to” means that one file, the source, contains a reference to another file, the target, rather than providing their own content. We found three types of references between Context Files:
@@
We found 518 file pairs that reference each other. The file type with most references was CLAUDE.md, with 357 “outgoing” references, followed by AGENTS.md and GEMINI.md. In contrast, the most frequently referenced file was AGENTS.md with 368 “incoming” references, substantially more than CLAUDE.md and copilot-instructions.md. Furthermore, the strongest pattern was the CLAUDE.md to AGENTS.md pair that occurred 311 times. In contrast, when AGENTS.md files acted as a pointer, they most frequently referenced CLAUDE.md files. This again reflects the symbiotic relationship between these two types of Context Files that we also found in the creation order discussed earlier.
@@ 6.2. Configuration Mechanism: Skills
Skills were introduced by Anthropic and now serve as an open standard for extending the capabilities of agentic AI tools with specialized contextual knowledge and workflows. Skills bundle prompts, tools, and documentation that an agent can invoke on demand. From an implementation perspective, Skills are a directory that contains a SKILL.md file. This file must include a YAML frontmatter with at least a name and description, followed by instructions in the body for how a task should be performed.
@@
As with Context Files, we also explored whether there are programming language-specific patterns. Figure 3 shows that TypeScript and Python are the dominating programming languages in all repositories that use agentic tools. This distribution is also reflected across projects that use Skills.
The specification of Skills recommends that SKILL.md files should contain fewer than 500 lines. It further recommends moving any detailed material to separate files (agentskills.io, 2026). We found that of the 601 Skills we identified, only 25 (4%) were longer than 500 lines. That is, most developers follow the recommendation. To provide detailed information and instructions, Skills can utilize different resources, i.e., additional optional files or directories that are loaded by the agents of a tool when required. There are three types of resource:
@@
To obtain the resources used by the different Skills, we scanned all Skills directories in the studied repositories for the occurrence of the relevant resource folders. Of the 601 skills analyzed, the vast majority (501, 83.3%) included no additional resources. Among those that did, the most common pattern was a references/ directory (48 skills, 8.0%), followed by a scripts/ directory (34, 5.7%). Twelve Skills (2.0%) combined both scripts/ and references/, while an assets/ directory appeared in only four Skills (0.7%). The remaining combinations, i.e., references/ with assets/, and all three directories together, each occurred just once (0.2%). In summary, Skills mostly use “static” resources (documentation that the agent reads when needed) rather than “dynamic” resources, such as executable scripts that extend agent behavior.
@@ 6.3. Configuration Mechanism: Subagents
Subagents are specialized AI agents that agentic AI coding tools or other agents can delegate tasks to. Technically, Subagents are a Markdown file with the same YAML frontmatter structure as Skills. However, Subagents differ in that they operate within their own context window and return the results to the parent agent whereas Skills execute instructions within the calling agent’s context (Anthropic, 2026; Cursor, 2026).
We found 452 Subagents across 131 repositories. On average, these repositories use 3.45 Subagents ($min=1.0$; $max=18.0$; $median=2$). However, Figure 5 indicates that these repositories commonly use fewer than three Subagents. One feature of Claude Code’s Subagents is that they can have their own “memory”, i.e., a persistent directory that survives across interactions with the agent (Anthropic, 2026). This directory can be used to build up knowledge over time, e.g., to store debugging insights. However, we could not find any repositories storing such memory files.
@@
  • Subagents: The usage patterns for Subagents are similar to those of Skills. Future work should study differences in content and use cases in more detail.
@@ 6.4. Threats to Validity
Our heuristics (Table 1) detect the presence of configuration artifacts but not whether the corresponding tool is actively used. We partially mitigate this by restricting our sample to actively maintained repositories, but artifact presence remains a proxy for adoption. Additionally, for GitHub Copilot, Cursor, and Gemini, detected files apply to both conversational and agentic modes, so we cannot isolate agentic usage. This limitation does not affect agentic-only mechanisms such as Skills and Subagents, nor the broader AGENTS.md trend.
@@ 7. Discussion
Our study offers a first cross-tool snapshot of configuration mechanisms and their artifacts across five agentic AI coding tools and 2,926 repositories. Below, we discuss the implications of our study.
@@
Our results reveal a clear trajectory toward AGENTS.md as a unifying, tool-agnostic configuration artifact. The creation order and reference patterns point in the same direction: For example, when both files are present, CLAUDE.md typically appears first and AGENTS.md is added later (Figure 7). CLAUDE.md most frequently points to AGENTS.md (311 cases), and AGENTS.md receives 368 incoming references overall, far more than any other file type. Repositories that began with copilot-instructions.md often later introduced CLAUDE.md or AGENTS.md, despite GitHub Copilot already supporting all major Context File types. Together, these patterns suggest bottom-up convergence around AGENTS.md as a cross-tool baseline, driven by developer practice and compatibility needs rather than vendor mandate. At the same time, the layering of multiple context files within a single repository raises the risk of redundant or conflicting instructions across artifacts. For tool vendors, native support for AGENTS.md—as already provided by Cursor and Codex—may become a baseline expectation.
@@
Across all three configuration mechanisms we examined in detail, adoption remains limited. Most repositories include only one or two Context Files (Figure 5), and the median number of Skills and Subagents per repository is two in both cases. Moreover, 83.3% of Skills do not include additional resources, and when resources are present, static documentation (references/) is more common than executable scripts (scripts/). In practice, Skills, therefore, function primarily as structured text rather than executable workflow bundles. This suggests that configuration is currently used more as documentation than as automation. We also found no evidence of repositories using the persistent memory feature of Claude Code’s Subagents.
This gap between available depth and actual usage likely reflects both the novelty of these mechanisms and the effort required to configure them. Claude Code was released only in February 2025, with Skills and Subagents introduced even more recently. Beyond recency, developers may gravitate toward the lowest-friction mechanism (i.e., Context Files) without exploring more advanced options. Defining executable Skills with scripts and structured resources requires additional design and maintenance effort compared to authoring Markdown instructions, which may deter adoption in the absence of clear guidance on best practices. At present, there is little empirical evidence on which configuration strategies are most effective or under which conditions they yield measurable improvements. Future work should therefore assess whether deeper configuration leads to performance gains, extending early evidence on the impact of Context Files (Lulla et al., 2026).
@@
Configuration practices vary systematically across tools. Claude Code repositories exhibit the broadest configuration footprint, whereas Cursor projects emphasize Rules and Commands, and Copilot and Codex repositories rarely extend beyond Context Files. These differences likely reflect both the configuration options each tool exposes (Table 1) and the norms emerging within their respective user communities. Tool ecosystems also differ in repository characteristics: Cursor repositories tend to be younger and larger, while Gemini repositories show comparatively higher activity levels (Table 3). An open question is whether these tool-specific configuration cultures will converge as feature sets overlap, or whether distinct patterns of use will persist.
@@
Our systematically compiled overview of eight configuration mechanisms (Table 1) provides a reference for developers seeking to understand the options available to them. Based on our findings, we offer the following concrete observations for practitioners. First, Context Files—particularly AGENTS.md—represent the lowest-friction entry point for configuring agentic tools and are already widely adopted. Second, developers who rely on multiple tools should consider maintaining an AGENTS.md file as a shared configuration baseline, given its cross-tool support and the reference patterns we observed. When multiple context files coexist, teams may benefit from clearly structuring them hierarchically (e.g., using tool-specific files as adapters that reference a shared core file) to reflect the layering patterns we identified. Third, for recurring, well-defined workflows, Skills offer the potential for richer configuration through scripts and structured resources, but this potential is currently underutilized. In practice, configuration is presently used more as documentation than as automation, suggesting that teams should adopt executable Skills deliberately and only when the expected benefits justify the additional setup effort. Finally, given the distinct configuration profiles across tools, teams should be mindful that adopting tool-specific mechanisms (e.g., rule-based configurations) may shape their workflows in tool-dependent ways.
For researchers, our findings point to several directions. Longitudinal studies are needed to track how configuration practices evolve as tools mature and as developers gain experience. Controlled studies should assess whether advanced mechanisms such as Skills with executable resources or dedicated Subagents provide measurable benefits over Context Files alone, extending prior work on the operational impact of AGENTS.md (Lulla et al., 2026). Given that projects increasingly use multiple tools (Section 5.1), research should also investigate interoperability challenges: how configuration artifacts interact across tools, and how potential conflicts between overlapping or contradictory instructions in different files can be detected and resolved. Our automated data collection and analysis pipeline (Authors, 2026) supports such ongoing analyses.
Finally, our study is limited to a single point-in-time snapshot of a rapidly evolving landscape. Agentic AI coding tools are under active development, and both the available configuration mechanisms and the patterns of their adoption are likely to change. The trends we identify—toward standardization around AGENTS.md, shallow adoption of advanced mechanisms, and tool-specific configuration cultures—are early empirical signals rather than settled findings.

### v2

@@ 5.1. Characterization of Repositories in Dataset
Before analyzing adoption, we characterize the 2,923 repositories in our dataset to contextualize the subsequent analysis.
Of the 2,923 repositories, 2,018 (69.0%) adopted a single tool, while 299 (10.2%) configured exactly two tools and 50 (1.7%) configured three or more. An additional 528 repositories (18.1%) use AGENTS.md as a tool-agnostic standard without any tool-specific configuration artifact; these are shown separately in the analysis as “AGENTS.md” (see Table 2, footnote d). Among repositories with multiple tools, Claude appeared most frequently with others. The most common combination was Claude and Copilot ($n=168$), followed by Claude and Cursor ($n=145$), Claude and Gemini ($n=54$), and Copilot and Cursor ($n=49$). Notably, 44.3% of Cursor repositories also configure Claude. The dominance of single-tool repositories limits the potential for multi-tool confounding in tool-specific analyses. In our initial sample, the top five main programming languages per repository ($n=36,184$) were Python (8,133; 22.5%); TypeScript (4,999; 13.8%); Java (3,813; 10.5%); Go (3,786; 10.5%); and JavaScript (3,709; 10.3%). Interestingly, this order is slightly different for repositories that use agentic coding tools ($n=2,923$, see Figure 2) where TypeScript and Go were more prominent and C# replaces JavaScript in the top five: TypeScript (745, 25.5%); Python (485, 16.6%); Go (457, 15.6%); C# (228, 7.8%); Java (227, 7.8%). TypeScript is the most common primary language for repositories using Claude, Copilot, and Cursor. Usage in Java and C# repositories is lower for all tools but Copilot. As shown in Table 3, which compares each tool’s adopting repositories against the complement of non-adopting repositories, repositories using Cursor are younger compared to those using other tools. Excluding Codex (nn=4), repositories associated with Gemini exhibit larger contributor counts and commit volumes, compared to those associated with other tools. Cursor repositories are the largest in terms of kilobytes of source code (excluding Codex). AGENTS.md repositories (those using only AGENTS.md with no tool-specific configuration) tend to have fewer commits and smaller size than the overall population, suggesting that these repositories adopted AGENTS.md as a lightweight, tool-agnostic starting point.
@@ 5.2. Distribution of Configuration Mechanisms
Figure 3 shows the distribution of configuration mechanisms per tool. Context Files (e.g., CLAUDE.md, AGENTS.md) were the most frequently adopted mechanism, used by 61.5–100% of repositories across all tools. Beyond Context Files, distinct tool-specific profiles emerge: 72.8% of Cursor repositories adopt Rules, since Cursor was one of the first to introduce this mechanism (Jiang and Nam, 2026), and 61.2% of Gemini repositories use Settings. Other mechanisms remain below 20% of repositories for most tools.
Most repositories included only a single Context File artifact, and the adoption of multiple non-context file mechanisms within the same repository remained rare. These results suggest that configuration usage is characterized by one dominant baseline configuration (i.e., Context Files).
Figure 4. Cumulative adoption of selected configuration artifacts for agentic tools. Copilot Instructions comprise two Context File artifact types (see Table 1).
Diagram showing the cumulative adoption of configuration artifacts for agentic tools.
@@
Diagram showing the configuration mechanism count per repository.
@@ 5.2.1. Co-occurrence of Configuration Mechanisms
We quantified pairwise associations between mechanism adoption using Cramér’s $V$ from chi-squared tests (28 comparisons, Benjamini–Hochberg corrected). Several configuration mechanisms are frequently adopted together, and some of these associations can be explained by shared configuration artifacts. For example, Settings and Hooks show a positive association ($V=0.36$), because hooks are usually defined in Settings files. Settings and Hooks also both co-occur with Skills ($V=0.15$ and $V=0.24$, respectively), which are defined and configured differently. Subagents also show positive associations with multiple mechanisms, including Settings, Skills, and MCP. Other configuration mechanisms appear less frequently in combination. Context Files are negatively associated with Rules ($V=0.41$) and Commands ($V=0.07$), which can be attributed to how they are defined and which tools first introduced the mechanisms. Overall, configuration mechanism adoption is characterized by recurring combinations rather than uniform usage, shaped by the tools and mechanisms they support (e.g., Claude, the dominating tool, does not support Rules). The full co-occurrence matrix is available in the supplementary material.
@@ 5.2.2. Adoption of Configuration Mechanisms over Time
Figure 4 shows the adoption over time of individual Context File artifacts, plus Skills and Subagents.
Context Files (e.g., CLAUDE.md, AGENTS.md) clearly dominated and increased continuously, while Skills and Subagents experienced comparatively slow growth. .cursorrules and copilot-instructions.md started being introduced in 2024, although Cursor and Copilot were originally released in 2023 and 2021, respectively; their agentic capabilities were only introduced in 2025 (see Table 2). The adoption of copilot-instructions.md has increased since then.
@@
RQ2 (Summary):
  • Context Files are the dominant and often sole configuration mechanism; other mechanisms are adopted less frequently. Over time, the gap has widened, with Skills and Subagents experiencing comparatively slow growth.
  • Adoption varies: Claude has a broad configuration footprint, while Cursor emphasizes rule-based mechanisms. Co-occurrence patterns are shaped by tool-specific mechanism support.
  • Most repositories (69.0%) adopt a single tool, and multi-tool overlap is limited. An additional 18.1% use AGENTS.md without any tool-specific configuration.
@@ 6. Details of Configuration Mechanisms (RQ3)
We analyzed three mechanisms in detail: Context Files and Skills because they are supported by all tools, and Subagents because they follow a similar format as Skills.
@@ 6.1. Configuration Mechanism: Context Files
Context Files are Markdown files that provide central, machine-readable context about a project. AGENTS.md, introduced by OpenAI, now serves as an open tool-agnostic convention increasingly supported by various tools (Mohsenimofidi et al., 2026). We identified 4,854 Context Files across 2,631 of the 2,923 repositories in our sample. Most repositories include one or two such files (Figure 5).
CLAUDE.md emerges as the dominant file type with 1,658 (34.2%), followed closely by AGENTS.md and copilot-instructions.md with 1,571 (32.4%) and 1,393 (28.7%) files, respectively. GEMINI.md (159 files, 3.3%) and .cursorrules (73 files, 1.5%) are rare. Note that .cursorrules are now deprecated; Cursor suggests using AGENTS.md instead. Cursor rules are different and are stored in .cursor/rules/. We also investigated the repository-level adoption of the Context Files: among the 2,631 repositories with Context Files, CLAUDE.md demonstrates the highest adoption rate with 45.4% of repositories (1,195 repos), AGENTS.md follows closely with 40.6% adoption (1,068 repos), and copilot-instructions.md appears in 35.4% of repositories (930 repos). Examining language-specific patterns, Context Files appeared most frequently in TypeScript (651 repos), Python (435), Go (414), C# (210), Java (209), JavaScript (200), Rust (180), C++ (156), PHP (123), and C (53) repositories. CLAUDE.md was the most common Context File across languages, except for Java, C#, and C++, where copilot-instructions.md dominated.
@@
Diagram showing the creation order of Context Files per repository.
@@
Diagram showing the reference patterns between Context Files.
Figure 6 shows the creation order of Context Files in repositories with multiple types (single-type repositories are omitted). CLAUDE.md was typically created first, with AGENTS.md commonly added afterward—likely because Claude Code is the most popular agentic tool but does not yet support the emerging standard AGENTS.md (anthropics/claude-code on GitHub, 2026). Repositories that started with copilot-instructions.md often later added CLAUDE.md or AGENTS.md, despite Copilot supporting all major Context File types, reinforcing their de facto standard status.
Finally, Context Files commonly reference each other (Figure 7). We identified three types of references, where a source file points to a target file rather than providing its own content:
@@
We found 518 reference pairs. CLAUDE.md had the most outgoing references (357), predominantly pointing to AGENTS.md (311 times). Conversely, AGENTS.md received the most incoming references (368) and, when acting as a pointer, most frequently referenced CLAUDE.md—reflecting the symbiotic relationship also observed in the creation order.
@@ 6.2. Configuration Mechanism: Skills
Skills were introduced by Anthropic and now serve as an open standard for extending agentic AI tools with specialized contextual knowledge and workflows. Skills bundle prompts, tools, and documentation that an agent can invoke on demand. Each Skill is a directory containing a SKILL.md file with YAML frontmatter (name, description) and instructions in the body.
@@
As with Context Files, TypeScript and Python dominated the language distribution of repositories that use Skills, mirroring their dominance across all repositories that use agentic tools (Figure 2).
The specification recommends that SKILL.md files contain fewer than 500 lines, with detailed material moved to separate files (agentskills.io, 2026). Only 29 of the 601 Skills (5%) exceeded this limit. To provide detailed information and instructions, Skills can include additional resource directories loaded by agents when required:
@@
We scanned all Skills directories for the occurrence of resource folders. Of the 601 skills analyzed, the vast majority (514, 85.5%) included no additional resources. Among those that did, the most common patterns were a references/ directory and a scripts/ directory (35 skills each, 5.8%). Eleven Skills (1.8%) combined both scripts/ and references/, while an assets/ directory appeared in only four Skills (0.7%). The remaining combinations, i.e., references/ with assets/, and all three directories together, each occurred just once (0.2%). In summary, Skills mostly use “static” resources (documentation that the agent reads when needed) rather than “dynamic” resources, such as executable scripts that extend agent behavior.
@@ 6.3. Configuration Mechanism: Subagents
Subagents are specialized AI agents to which agentic AI coding tools or other agents can delegate tasks. They share the same YAML frontmatter structure as Skills but operate within their own context window and return results to the parent agent, whereas Skills execute within the calling agent’s context (Anthropic, 2026; Cursor, 2026).
We found 450 Subagents across 131 repositories. On average, these repositories use 3.44 Subagents ($min=1.0$; $max=17.0$; $median=2$), though most use fewer than three (Figure 5). Claude Code’s Subagents can have their own “memory,” i.e., a persistent directory that survives across interactions and can build up knowledge over time, e.g., to store debugging insights (Anthropic, 2026). However, we found no repositories storing such memory files.
@@
  • Subagents: Usage patterns mirror those of Skills, with most repositories defining only one or two. No repositories use the persistent memory feature available for Claude Code’s Subagents.
@@ 6.4. Threats to Validity
Our heuristics (Table 1) detect the presence of configuration artifacts but not whether the corresponding tool is actively used. We partially mitigate this by restricting our sample to actively maintained repositories, but artifact presence remains a proxy for adoption. As a further mitigation, we scanned the complete commit history of every repository with at least one configuration file for AI-authored commits, using author/committer identity fields, git trailers (e.g., Co-authored-by), and commit message patterns. Among the 2,895 repositories with at least one configuration artifact, 2,079 (71.8%) also contain AI-authored commits, providing evidence that detected artifacts are associated with active tool usage rather than being dormant files. The detection scripts and full results are included in the supplementary material. Additionally, for GitHub Copilot, Cursor, and Gemini, detected files apply to both conversational and agentic modes, so we cannot isolate agentic usage. This limitation does not affect agentic-only mechanisms such as Skills and Subagents, nor the broader AGENTS.md trend. Although multi-tool overlap could conflate tool-specific patterns, 85.3% of tool-adopting repositories configure a single tool, limiting this concern. To verify that detected artifacts contain meaningful content, we programmatically scanned all configuration files across the eight artifact types and excluded empty files (reducing the initial 2,926 detected repositories to 2,923). One of the authors then used Claude Code (Opus 4.6 with high effort) to check all 451 files with ten or fewer lines. Of these, 396 contained substantive configuration content (e.g., coding conventions, agent definitions, tool settings), 51 were reference-only files analyzed separately in Section 6, and four were borderline cases (e.g., a bare heading or filename). The supplementary material includes the inspection script, the checked files, and their classification.
@@ 7. Discussion
Our study offers a first cross-tool snapshot of configuration mechanisms and their artifacts across five agentic AI coding tools and 2,923 repositories. Below, we discuss the implications of our study.
@@
Our results reveal a clear trajectory toward AGENTS.md as a unifying, tool-agnostic configuration artifact. Creation order analysis shows that CLAUDE.md typically appears first with AGENTS.md added later (Figure 6). Reference patterns confirm this. CLAUDE.md most frequently points to AGENTS.md (311 cases), which receives 368 incoming references overall, far more than any other file type. Together, these patterns suggest bottom-up convergence around AGENTS.md as a cross-tool baseline, driven by developer practice and compatibility needs rather than vendor mandate. This is further reinforced by the 528 repositories that use AGENTS.md without any tool-specific configuration, suggesting that developers adopt it as a tool-agnostic standard even before committing to a specific tool. At the same time, the layering of multiple context files within a single repository raises the risk of redundant or conflicting instructions across artifacts. For tool vendors, native support for AGENTS.md—as already provided by Cursor and Codex—may become a baseline expectation.
@@
Across all three configuration mechanisms we examined in detail, adoption remains limited. Most repositories include only one or two Context Files (Figure 5), and the median number of Skills and Subagents per repository is two in both cases. Moreover, 85.5% of Skills do not include additional resources, and when resources are present, static documentation (references/) is equally common as executable scripts (scripts/). In practice, Skills, therefore, function primarily as structured text rather than executable workflow bundles. This suggests that configuration is currently used more as documentation than as automation. We also found no evidence of repositories using the persistent memory feature of Claude Code’s Subagents.
This gap likely reflects both the novelty of these mechanisms and the effort required to configure them. Developers may gravitate toward the lowest-friction mechanism (i.e., Context Files), since defining executable Skills with scripts and structured resources requires more design and maintenance effort than authoring generic Markdown-based files. Future work should assess whether deeper configuration leads to measurable performance gains, extending early evidence on the impact of Context Files (Lulla et al., 2026).
@@
Configuration practices vary systematically across tools: Claude Code repositories exhibit the broadest configuration footprint, Cursor projects emphasize Rules and Commands, and Copilot repositories rarely extend beyond Context Files. These differences likely reflect both the configuration options each tool exposes (Table 1) and the norms emerging within their respective user communities. Repository characteristics also differ: Cursor repositories tend to be younger and larger, while Gemini repositories show higher activity levels (Table 3). The 528 repositories using AGENTS.md without any tool-specific configuration further reinforce its role as a tool-agnostic standard. It remains an open question whether these tool-specific configuration cultures will converge as tool capabilities increasingly overlap, or whether distinct usage patterns will persist.
@@
Table 1 provides a reference for developers seeking to understand available configuration options. Based on our findings, we offer the following observations for practitioners. First, Context Files—particularly AGENTS.md—represent the lowest-friction entry point for configuring agentic tools and are already widely adopted. Second, developers who rely on multiple tools should maintain an AGENTS.md file as a shared configuration baseline, given its cross-tool support and the reference patterns we observed. When multiple context files coexist, teams may benefit from structuring them hierarchically (e.g., using tool-specific files as adapters that reference a shared core file). Third, for recurring, well-defined workflows, Skills offer potential for richer configuration through scripts and structured resources, but in practice Skills are presently used more as static documentation than as executable workflows. Finally, adopting tool-specific mechanisms (e.g., rule-based configurations) may shape workflows in tool-dependent ways.
For researchers, our findings point to several directions. Longitudinal studies are needed to track how configuration practices evolve as tools mature. Controlled studies should assess whether advanced mechanisms such as Skills with executable resources or dedicated Subagents provide measurable benefits over Context Files alone (Lulla et al., 2026). Given that certain projects use multiple tools (Section 5.1), research should also investigate interoperability challenges, including how conflicts between overlapping or contradictory instructions across files can be detected and resolved. Our automated data collection and analysis pipeline (Authors, 2026) supports such ongoing analyses.
Finally, our findings represent a point-in-time snapshot of a rapidly evolving landscape. The trends we identify—toward standardization around AGENTS.md, shallow adoption of advanced mechanisms, and tool-specific configuration cultures—are early empirical signals rather than settled findings.